OpenStudy (anonymous):

Joe traveled to the town hall and back. It took 3 hours long to go there than it did to come back. The average speed on the trip was 30km/h. The average speed on the way back was 75km/h. How many hours did the trip there take?

hero (hero):

d = rt To get there, it takes three hours longer so in other words, three hours slower: d1 = 30(t - 3) Coming back, it is three hours quicker: d2 = 75t No matter how slow or fast the trip, the distance to and from two two specific locations will always be the same, so: d1 = d2 30(t - 3) = 75t Solve for t

OpenStudy (anonymous):

Thank you!!

hero (hero):

Did you solve for t yet?

OpenStudy (anonymous):

So far i got 30t-90= 75t then 30 and 75 which would be 135t................. and i think i did it wrong

hero (hero):

Actually, after further investigation, the formula should be 30(t + 3) = 75t

OpenStudy (anonymous):

uh explain please

hero (hero):

Well, basically, t represents how long, in hours, it will take to travel to get from A to B. Since getting there takes longer, it will take t + 3 hours.

OpenStudy (anonymous):

So then wouldn't you distribute the 30 to the T and the +3